# ”Living Within Four Walls”: Exploring Emotional and Social Dynamics in Mobile Usage During Home Confinement

### v3

@@ 5.3.1. General Model
Table 6. Prediction performance of the general models for valence and arousal with different feature sets. The regression models are K-Nearest Neighbor (KNN), Random Forest (RF) and Gradient Boosting (GB), compared with Mean and Median baseline. The evaluation metrics are RMSE and MAE (with standard deviation).
Prediction Result (RMSE(SD)/MAE(SD))
Mean Baseline Median Baseline KNN Regressor RF Regressor GB Regressor
@@
[…]
Features with $F$-statistic ¿= 5
[…]
Features with $F$-statistic ¿= 10
[…]
@@
5.3.2. Group Models
We gathered participant work start times from pre-study questionnaires and categorized participants into two distinct groups: $EARLY$ Workers (678 data points from 12 participants who typically commence work between 7:00 and 10:00 AM) and $LATE$ Workers (778 data points from 13 participants who begin work after 3:00 PM). For the early workers, the average self-reported $valence$ is 3.2 (±$\pm 0.9$) and $arousal$ is 2.9 (±$\pm 1.0$). For the late workers, the average $valence$ is 3.7 (±$\pm 1.0$) and $arousal$ is 2.6 (±$\pm 1.2$). Prediction models were subsequently developed for each group using all features.
@@ 5.3.2. Group Models
Figure 7. Prediction performance of $valence$ and $arousal$ models for two user groups: $EARLY$ workers (typically starting work between 7:00 and 10:00 AM) and $LATE$ workers (starting work after 3:00 PM). The models are evaluated with 3-fold cross-validation, where the error bar presents the standard deviation.
@@ 5.3.3. Individual Models
(a) MAE for valence
(b) RMSE for valence
(c) MAE for arousal
(d) RMSE for arousal
Figure 8. Prediction performance of individual models using all features
5.4. Feature Importance
@@
Figure 9. Top 15 features based on impurity score with RF regressors of the general models.
As depicted in Figure 9, we observe that the most important features for both valence and arousal are similar, but with different weights. These features primarily include time-related, app-category, and notification features. Both general models consistently identify ‘sin_time’ and ‘cos_time’ as the top two important features, indicating that the time period of the day greatly influences participants’ valence and arousal. Additionally, the day of the week (‘weekday’) impacts both models, although it has less importance in the arousal model.
Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features in the general models, but it appears as important in each group models. It worth mentioning that ‘time_period’ is the most important feature for predicting $valence$ for $EARLY$ workers, but not for $LATE$ workers. These findings differ slightly from the observations reported in (35), which found that app usage patterns varied greatly by time and day. The discrepancy may be attributed to the unique circumstances of home confinement, where individuals may experience a loss of time perception and a blurring of the distinction between weekdays and weekends. This aligns with the findings presented in (12), which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
Figure 10. Top 5 features based on impurity score with RF regressors of group models.
Besides, as shown in Figure 10, for $LATE$ workers, the same five features with similar importance distributions are observed for both $valence$ and $arousal$ predictions, all are related to app interaction. In contrast, $EARLY$ workers show a more diverse feature set; while the interacted app notification is the second most important, physical activity also contribute to $valence$ prediction. This difference may reflect the broader range of daytime activities available to $EARLY$ workers, potentially explaining their lower prediction performance discussed in Section 5.3.2. For $LATE$ workers, the reliance on app interaction highlights a potential concern regarding phone dependency.
6. Discussion and Implications
This study explores the impact of home confinement during the COVID-19 lockdown on individuals’ well-being, social roles, and work-life balance, using unobtrusive mobile phone sensing to gain insights into participants’ emotional states and behaviors. Although the study specifically focuses on the COVID-19 lockdown, we argue that its findings have broader relevance, especially in light of the growing trend of remote work and other situations of enforced isolation.
6.1. Well-being and Social Roles During Home Confinement (RQ1)
@@
Although the findings are rooted in the context of the COVID-19 lockdown, the implications extend beyond this unique situation. Home confinement is not limited to pandemic scenarios but is increasingly relevant to other forms of isolation or reduced mobility, such as remote work, post-operative recovery, disability-related confinement, or even digital addiction. The growing trend of telework and flexible work arrangements in the post-pandemic world further underscores the importance of understanding the emotional and behavioral impacts of prolonged confinement in various forms.
6.2. Predicting Emotional States Through Mobile Usage (RQ2)
@@
6.3. Design Implications
Based on the findings of this study, we propose several design implications for future interventions and mobile applications aimed at supporting well-being during periods of home confinement or isolation:
Personalized Predictive Models. The success of time-related features in predicting emotional states suggests that mobile apps should incorporate personalized, time-sensitive data to offer real-time emotional tracking and tailored interventions. For example, apps could adjust their notifications and suggestions based on users’ routines, ensuring that interventions are timely and relevant.
Balanced Digital Interactions. Our study found that interruptions, particularly from communication apps, may exacerbate the blurring of work-life boundaries. Future apps should focus on helping users manage interruptions, offering customizable notification settings or “focus mode” features to reduce disruptions and help maintain better separation between work and personal life.
Encouraging Healthy Digital Habits. Tools that support productivity, such as task managers and calendars, were associated with higher valence. Future designs should encourage healthy digital habits by integrating features that promote structure, goal-setting, and task management. Additionally, incorporating reminders for physical activity or mental health breaks could further support users’ emotional well-being during periods of isolation.
Context-Aware Interventions. Given the differences in how users experience isolation based on their routines, digital interventions should be context-aware and adaptive. For example, users with later work hours may benefit from interventions encouraging physical activity or relaxation earlier in the day, while users with earlier start times may need more focused strategies for managing digital distractions during the day.
6.4. Generalizability of the Findings
Although this study specifically focuses on the COVID-19 home confinement context, we believe the findings can extend to a range of other isolation scenarios. While the emotional triggers—such as fear of infection in a pandemic or frustration with limited mobility in recovery—may differ, the core psychological mechanisms and emotional dynamics of isolation are broadly similar. Our study observed patterns like emotional detachment, stress, and blurred work-life boundaries, all of which are commonly experienced during prolonged isolation, regardless of the specific circumstances.
The emotional dynamics of isolation—such as changes in social interactions, work-related stress, and adjusting to new routines—are consistent across scenarios like remote work, post-surgical recovery, or other forms of confinement. For example, both remote workers and those recovering from surgery face challenges with work-life balance and social isolation. Our findings about app usage—especially productivity apps promoting positive emotional states—are relevant in these contexts, suggesting that digital tools can support well-being across various isolation scenarios.
Additionally, the mobile sensing approach used in this study, which tracks behaviors like app usage and interaction timing, can be applied beyond COVID-19. While the types of apps and behaviors may vary, the principle of using mobile devices to track emotional states in real-time is adaptable to different situations, such as remote work or recovery periods. The ability to monitor and intervene based on individual usage patterns could be valuable in a wide range of confinement contexts.
@@ 6.4. Generalizability of the Findings
7. Limitations
This study has several limitations that must be acknowledged, which may affect the validity of our findings (78):
@@ 7. Limitations
Device Usage and Untracked Data (Threats to Internal and Construct Validity). A key limitation is the potential for incomplete data due to participants using untracked devices, such as tablets or additional smartphones, or even desktops while only contributing smartphone data. These unmonitored interactions could have influenced emotional states and introduced noise into the data. The absence of comprehensive tracking across all devices limits our ability to fully assess participants’ total digital behavior and emotional responses, suggesting the need for more extensive device tracking in future studies.
@@
8. Conclusion
In this field study, we explored how human emotions, social roles, and mobile usage behaviours were affected during the COVID-19 home confinement period. This rare event offers optimal conditions to while minimizing confounding factors like nature and physical social interactions. It contributes to understanding the mental health implications of heavy reliance on technology and work-from-home arrangements. By tracking mental well-being, the study highlights how individuals manage work-life balance and adapt their schedules or arrangements (e.g., work-from-home or half-office half-home) effectively. Additionally, it provides valuable insights into psychological states and smartphone usage behaviours during confinement, informing interventions for similar scenarios, such as post-operative home rehabilitation. Understanding how people navigate home confinement and the role of technology in their lives can guide the design of smart environments and remote support programs, improving care quality. These findings also have broader implications for understanding human behaviour during crises, supporting the development of policies and interventions for future challenges.
Despite the challenges we faced in collecting data during COVID-19 lockdowns, we believe that gaining insights into human well-being and mobile usage behaviours during home confinement is crucial. This understanding is not only valuable for policymakers in implementing effective measures during future outbreaks, but also for individuals to become more self-aware of their well-being and mental health. Moving forward, our research will expand to investigate users’ desktop usage behaviour. By comparing the differences in usage behaviour across different platforms, we aim to gain a deeper understanding of people’s daily lives and work-life balance.
9. Acknowledgments

### v4

@@ 5.3.1 General Model
Table 6: Prediction performance of the general models for valence and arousal with different feature sets. The regression models are K-Nearest Neighbor (KNN), Random Forest (RF) and Gradient Boosting (GB), compared with Mean and Median baseline. The evaluation metrics are RMSE and MAE (with standard deviation).
RMSE/MAE(SD) Mean Baseline Median Baseline KNN Regressor RF Regressor GB Regressor
@@
Valence (29) 0.986(.002)/0.824(.002) 1.091(.003)/0.767(.004) 0.963(.026)/0.776(.022) 0.903(.025)/0.717(.021) 0.925(.020)/0.743(.012)
Arousal (29) 1.113(.004)/0.884(.004) 1.146(.004)/0.784(.004) 1.088(.023)/0.844(.019) 1.005(.028)/0.766(.020) 1.084(.044)/0.756(.037)
Features with $F$-statistic $\geq 5$
Valence (17) 0.986(.002)/0.824(.002) 1.091(.003)/0.767(.004) 0.956(.020)/0.769(.018) 0.904(.024)/0.717(.021) 0.986(.020)/0.725(.021)
Arousal (14) 1.113(.004)/0.884(.004) 1.146(.004)/0.784(.004) 1.064(.027)/0.818(.026) 1.015(.024)/0.778(.018) 1.087(.032)/0.747(.022)
Features with $F$-statistic $\geq 10$
Valence (13) 0.986(.002)/0.824(.002) 1.091(.003)/0.767(.004) 0.936(.020)/0.730(.013) 0.912(.024)/0.725(.018) 0.987(.046)/0.724(.019)
Arousal (12) 1.113(.004)/0.884(.004) 1.146(.004)/0.784(.004) 1.064(.027)/0.819(.026) 1.011(.023)/0.774(.017) 1.076(.035)/0.745(.030)
@@
5.3.2 Group Models
We gathered participant work start times from pre-study questionnaires and categorized participants into two distinct groups: EARLY Workers (678 data instances from 12 participants who typically commence work between 7:00 and 10:00 AM) and LATE Workers (778 data instances from 13 participants who begin work after 3:00 PM). For the early workers, the average self-reported valence is 3.2 (±$\pm 0.9$) and arousal is 2.9 (±$\pm 1.0$). For the late workers, the average valence is 3.7 (±$\pm 1.0$) and arousal is 2.6 (±$\pm 1.2$). Prediction models were subsequently developed for each group using all features.
@@ 5.3.2 Group Models
Figure 7: Prediction performance of valence and arousal models for two user groups: EARLY workers (typically starting work between 7:00 and 10:00 AM) and LATE workers (starting work after 3:00 PM). The models are evaluated with 3-fold cross-validation, where the error bar presents the standard deviation.
@@ 5.3.3 Individual Models
(a) valence (MAE)
(b) valence (RMSE)
(c) arousal (MAE)
(d) arousal (RMSE)
Figure 8: Prediction performance of individual models using all features
5.4 Feature Importance
@@
Figure 9: Top 15 features based on impurity score with RF regressors of the general models.
As depicted in Figure 9, we observe that the most important features for both valence and arousal are similar, but with different weights. These features primarily include time-related, app-category, and notification features.
Both general models consistently identify ‘sin_time’ and ‘cos_time’ as the top two important features, indicating that the time period of the day greatly influences participants’ valence and arousal. Additionally, the day of the week (‘weekday’) impacts both models, although it has less importance in the arousal model.
Surprisingly, the feature ‘is_weekend’ does not appear among the top 15 features in the general models, but it appears as important in each group model. It is worth mentioning that ‘time_period’ is the most important feature for predicting valence for EARLY workers, but not for LATE workers. These findings differ slightly from the observations reported in [75], which found that app usage patterns varied greatly by time and day. The discrepancy may be attributed to the unique circumstances of home confinement, where individuals may experience a loss of time perception and a blurring of the distinction between weekdays and weekends. This aligns with the findings presented in [87], which suggest that the perception of time and daily routines can be disrupted during periods of confinement.
Figure 10: Top 5 features based on impurity score with RF regressors of group models.
Besides, as shown in Figure 10, for LATE workers, the same five features with similar importance distributions are observed for both valence and arousal predictions, all are related to app interaction. In contrast, EARLY workers show a more diverse feature set; while the interacted app notification is the second most important, physical activity also contributes to valence prediction. This difference may reflect the broader range of daytime activities available to EARLY workers, potentially explaining their lower prediction performance discussed in Section 5.3.2. For LATE workers, the reliance on app interaction highlights a potential concern regarding phone dependency.
6 Discussion and Implications
This study explores the impact of home confinement on individuals’ well-being, social roles, and work-life balance, using unobtrusive mobile phone sensing to gain insights into participants’ emotional states and behaviors. Although the study specifically focuses on the COVID-19 lockdown, we argue that its findings have broader relevance, especially in light of the growing trend of remote work and other situations of enforced isolation.
6.1 Well-being and Social Roles During Home Confinement (RQ1)
@@
In wider contexts, home confinement is not limited to pandemic scenarios but is increasingly relevant to other forms of isolation or reduced mobility, such as remote work, post-operative recovery, disability-related confinement, or even digital addiction. The growing trend of telework and flexible work arrangements in the post-pandemic world further underscores the importance of understanding the emotional and behavioral impacts of prolonged confinement in various forms.
6.2 Predicting Emotional States Through Mobile Usage (RQ2)
@@
6.3 Generalizability of the Findings
Although this study specifically focuses on the COVID-19 home confinement context, we believe the findings can extend to a range of other isolation scenarios. While the emotional triggers – such as fear of infection in a pandemic or frustration with limited mobility in recovery – may differ, the core psychological mechanisms and emotional dynamics of isolation are broadly similar. Our study observed patterns like emotional detachment, stress, and blurred work-life boundaries, all of which are commonly experienced during prolonged isolation, regardless of the specific circumstances.
The emotional dynamics of isolation – such as changes in social interactions, work-related stress, and adjusting to new routines – are consistent across scenarios like remote work, post-surgical recovery, or other forms of confinement. For example, both remote workers and those recovering from surgery face challenges with work-life balance and social isolation. Our findings about app usage—especially productivity apps promoting positive emotional states—are relevant in these contexts, suggesting that digital tools can support well-being across various isolation scenarios.
Additionally, the mobile sensing approach used in this study, which tracks behaviors like app usage and interaction timing, can be applied beyond the studied context. While the types of apps and behaviors may vary, the principle of using mobile devices to track emotional states in real-time is adaptable to different situations, such as remote work or recovery periods. The ability to monitor and intervene based on individual usage patterns could be valuable in a wide range of confinement contexts.
@@ 6.3 Generalizability of the Findings
7 Limitations
This study has several limitations that must be acknowledged, which may affect the validity of our findings [88]:
@@ 7 Limitations
Device Usage and Untracked Data (Threats to Internal and Construct Validity). A key limitation is the potential for incomplete data due to participants using untracked devices, such as tablets or additional smartphones, or even desktops, while only contributing smartphone data. These unmonitored interactions could have influenced emotional states and introduced noise into the data. The absence of comprehensive tracking across all devices limits our ability to fully assess participants’ total digital behavior and emotional responses, suggesting the need for more extensive device tracking in future studies.
@@
8 Conclusion
In this field study, we explored how human emotions, social roles, and mobile usage behaviors were affected during the home confinement period. This rare event offers optimal conditions while minimizing confounding factors like nature and physical social interactions. It contributes to understanding the mental health implications of heavy reliance on technology and work-from-home arrangements. By tracking mental well-being, the study highlights how individuals manage work-life balance and adapt their schedules or arrangements (e.g., work-from-home or half-office half-home) effectively. Additionally, it provides valuable insights into psychological states and smartphone usage behaviors during confinement, informing interventions for similar scenarios, such as post-operative home rehabilitation. Understanding how people navigate home confinement and the role of technology in their lives can guide the design of smart environments and remote support programs, improving care quality. These findings also have broader implications for understanding human behavior during crises, supporting the development of policies and interventions for future challenges.
Despite the challenges we faced in collecting data during the lockdowns, we believe that gaining insights into human well-being and mobile usage behaviors during home confinement is crucial. This understanding is not only valuable for policymakers in implementing effective measures during future outbreaks, but also for individuals to become more self-aware of their well-being and mental health. Moving forward, our research will expand to investigate users’ desktop usage behavior. By comparing the differences in usage behavior across different platforms, we aim to gain a deeper understanding of people’s daily lives and work-life balance.
9 Acknowledgments